# Rajiv Raman

### v2

@@ Abstract
All our results are algorithmic and yield polynomial time algorithms (when the treewidth is bounded). The existence and construction of sparse supports is a crucial step in the design and analysis of PTASs and/or sub-exponential time algorithms for several packing and covering problems.
@@ 1 Introduction
A hypergraph is defined by a set $X$ of vertices and a collection $\mathcal{H}$ of subsets of $X$, called hyperedges. A support for $(X,\mathcal{H})$ is a graph $Q=(X,F)$ s.t. $Q[H]$ is connected for each $H\in\mathcal{H}$, where $Q[H]$ is the subgraph of $Q$ induced by the vertices in $H$.
Voloshina and Feinberg introduced the notion of a support as a notion of planarity of hypergraphs [voloshina1984planarity]. The notion is now well studied in hypergraph visualization [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored] and network design [anceaume2006semantic, baldoni2007tera, baldoni2007efficient, chand2005semantic, hosoda2012approximability, korach2003clustering, onus2011minimum]. The notion of supports and similar notions have also found applicability for geometric problems [Mustafa17, BasuRoy2018, Cohen-AddadM15, krohn2014guarding, mustafa2010improved, RR18].
Every hypergraph trivially admits a support by putting a complete graph on the vertices. However, we are interested in the existence and construction of sparse supports. Johnson and Pollak [Johnson1987Hypergraph] showed that it is NP-complete to decide if a hypergraph admits a support that is a planar graph.
We consider hypergraphs defined by subgraphs of a graph. Let $G=(V,E)$ be a given graph and let $\mathbf{b}(V)$ and $\mathbf{r}(V)$ be a partition of $V(G)$ into a set of terminals and a set of non-terminals, respectively. A family $\mathcal{H}$ of subgraphs of $G$ defines a hypergraph $(\mathbf{b}(V),\mathcal{H})$, whose vertex set is $\mathbf{b}(V)$ and each $H\in\mathcal{H}$ defines a hyperedge consisting of the vertices in $\mathbf{b}(V)\cap V(H)$. We call this hypergraph the primal hypergraph. Similarly, we define the dual hypergraph whose elements are the subgraphs in $\mathcal{H}$ and each $v\in\mathbf{b}(V)$ defines a hyperedge consisting of all $H\in\mathcal{H}$ containing $v$. We also consider a generalization of the primal and dual hypergraphs namely, the intersection hypergraphs. Let $\mathcal{H}$ and $\mathcal{K}$ be two families of subgraphs of $G$. The intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ is the hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}}),\mbox{ where }\mathcal{H}_{K}=\{H\in\mathcal{H}:V(H)\cap V(K)\neq\emptyset\}$
@@
A support for a primal (resp. dual/intersection) hypergraph is called a primal (resp. dual/intersection) support. We design algorithms to construct primal, dual and intersection supports for graphs of bounded treewidth. We obtain similar results for a special case of graphs of treewidth 2, namely, the outerplanar graphs.
@@
In order to construct an intersection support in the case of bounded treewidth graphs, we require the construction of primal and dual supports. In the dual setting, note that we can assume without loss of generality that $\mathbf{b}(V)=V$.
@@
As we show below, for the problem defined above, the answer is negative even if $\mathcal{G}$ is the family of trees. However, if we restrict the family of subgraphs to be non-piercing, then the answer to the question above is affirmative.
Definition 1 (Non-piercing).
The notion of non-piercing is motivated by the results of Raman and Ray [RR18], where the authors gave algorithms to construct planar primal, dual and intersection supports for hypergraphs defined by non-piercing regions [^2] A collection $\mathcal{R}$ of connected regions in general position in the plane is called non-piercing if, for each $R,R^{\prime}\in\mathcal{R}$, the region $R\setminus R^{\prime}$ is a connected region in the plane. in the plane. Recently, Raman and Singh [raman2025supportsgraphsboundedgenus] generalized the results of [RR18] to show that if subgraphs of a graph embedded on an oriented surface are cross-free, then there is a support also of the same genus as that of the host graph. Their results imply a unified analysis for several packing and covering problems defined on the hypergraphs. In particular, the existence of a sparse support in many cases implies a PTAS for these problems via a framework of local-search algorithms. See [DBLP:conf/walcom/AschnerKMY13, ChanH12, mustafa2010improved] for a description of the local-search framework. We give two candidate examples where our techniques apply.
Consider the following problem: A collection of cliques in a graph is clearly a collection of non-piercing subgraphs. Given a graph $G$ and a parameter $k\in\mathbb{N}$, find a smallest subset of vertices that hits all cliques of size at least $k$, or a more general version of this problem - Given a set $\mathcal{H}$ of cliques and a set $\mathcal{K}$ of cliques, find a smallest subset of cliques in $\mathcal{H}$ that hit all the cliques in $\mathcal{K}$. As a second example, let $G$ be a graph and let $\mathcal{H}$ be a collection of non-crossing paths in $G$ where we say that two paths $P_{1}$ and $P_{2}$ are non-crossing if once the paths meet at a vertex, they do not separate. Non-crossing paths are, by definition, non-piercing. Suppose we want to select the smallest number of edge-disjoint, or vertex-disjoint paths from $\mathcal{H}$. If $G$ has bounded treewidth, then our results imply the existence of a support of bounded treewidth for the two examples above. Hence, the two problems described above admit a PTAS via a local-search framework.
@@
The non-piercing, or cross-free conditions were used to obtain supports for geometric regions in the plane, or for graphs of bounded genus [raman2025supportsgraphsboundedgenus]. In this paper, since the graph classes are simple, one may wonder if the non-piercing condition is necessary to obtain sparse supports. The following examples show that this is indeed the case. For the primal hypergraph, consider a star $K_{1,n}$ with leaves $v_{1},\ldots,v_{n}$, and $v_{0}$, the central vertex. We color the leaves $\mathbf{b}$ and the central vertex $\mathbf{r}$. The subgraphs $\mathcal{H}$ consist of all pairs of leaves and the central vertex $v_{0}$. Thus, =$\mathcal{H}=\{H_{ij}:H_{ij}=\{v_{i},v_{0},v_{j}\}\}$. It is easy to see that the primal support in this case is $K_{n}$, the complete graph on $n$ vertices. For the dual hypergraph, consider a star $K_{1,\binom{n}{2}}$ with central vertex $v_{0}$. Each leaf $v_{j}$ is labeled by a unique pair $\{x_{j},y_{j}\}$ of $\{1,\ldots,n\}$. There are $n$ subgraphs $H_{1},H_{2},\ldots,H_{n}$, where $H_{i}=\{v_{0}\}\cup\{v_{j}:i\in\{x_{j},y_{j}\}\}$. It is easy to check that the dual support in this case is also $K_{n}$, as there are exactly two subgraphs containing each leaf.
While the notion of supports has found applications in different domains, we still do not have a good tool-set for deciding or constructing supports even in restricted settings, many of which arise in applications. For example, Buchin et al. [buchin2011planar] showed that deciding if a hypergraph admits a 2-outerplanar [^3] A graph $G$ is 2-outerplanar if there is an embedding of $G$ in the plane such that removing the vertices of the outer face results in an outerplanar graph. support is NP-hard, and they left open the question of deciding if a hypergraph admits an outerplanar support. While we do not answer their question, we believe that the tools we develop in constructing an outerplanar support may help in eventually resolving their question.
@@ 2 Related Work
The notion of planarity for hypergraphs, unlike that for graphs is not uniquely defined. This notion was first studied by Zykov [zykov], but that was very restrictive. Voloshina and Feinberg [voloshina1984planarity] defined a notion of planarity of hypergraphs that is equivalent to the existence of a support graph that is planar, i.e., a planar support. Johnson and Pollak [Johnson1987Hypergraph] showed that deciding if a hypergraph admits a planar support is NP-hard. Buchin et al. [buchin2011planar] strengthened their result showing that it is NP-hard even to decide if a hypergraph admits a support that is $k$-outerplanar, for $k\geq 2$. However, one can decide in polynomial time if a hypergraph has a support that is a path, a cycle, a cactus [^4] A connected graph is called cactus if each of its edges participates in at most one cycle., or a tree with bounded maximum degree [BOOTH1976335, brandes2010blocks, buchin2011planar, tarjan1984simple].
@@ 3 Preliminaries
Let $\mathbf{b}(V)$ and $\mathbf{r}(V)$ be a partition of $V(G)$ into a set of terminals and a set of non-terminals, respectively. Formally, let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be a 2-coloring of $V(G)$ (not necessarily proper). We call $\mathbf{r}(V)=c^{-1}(\mathbf{r})$ the set of red vertices and $\mathbf{b}(V)=c^{-1}(\mathbf{b})$ the set of blue vertices. We use $\mathbf{b}(H)$ to denote $V(H)\cap\mathbf{b}(V)$, and $\mathbf{r}(H)$ to denote $V(H)\cap\mathbf{r}(V)$. Since two subgraphs are said to intersect if they share a vertex, a hypergraph defined by subgraphs of a graph $G$ remains the same if the subgraphs are replaced by the corresponding induced subgraphs. Moreover, if $\mathcal{H}$ is a collection of non-piercing subgraphs, then it remains non-piercing if the subgraphs are replaced by the corresponding induced subgraphs. Therefore, we assume throughout that the term subgraph refers to an induced subgraph of $G$.
@@
For a graph $G$ and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs of $G$, we use the pair $(G,\mathcal{H})$ to refer to the primal or dual hypergraph, and triple $(G,\mathcal{H},\mathcal{K})$ for the intersection hypergraph. We call the tuple $(G,\mathcal{H})$ a graph system and the tuple $(G,\mathcal{H},\mathcal{K})$ an intersection system. If $G\in\mathcal{G}$, and $\mathcal{H,K}$ satisfy a property $\Pi$, then we say that $(G,\mathcal{H})$ is a $\mathcal{G}$ $\Pi$ graph system, or that $(G,\mathcal{H},\mathcal{K})$ is a $\mathcal{G}$ $\Pi$ intersection system. For example, if $G$ has bounded treewidth, and $\mathcal{H}$ and $\mathcal{K}$ are connected subgraphs of $G$, then $(G,\mathcal{H},\mathcal{K})$ is a bounded treewidth connected intersection system. If $G$ is outerplanar and $\mathcal{H}$ is non-piercing, then $(G,\mathcal{H})$ is an outerplanar non-piercing system.
Note that adding edges to $G$ leaves the subgraphs in $\mathcal{H}$ or $\mathcal{K}$ non-piercing. Thus, we can assume in the outerplanar setting that $G$ is a maximal outerplanar graph, and in the setting where $G$ has bounded treewidth, we can assume that each bag in a tree decomposition of the graph induces a complete graph, as this does not increase the treewidth (See Section 6 for more details).
@@
We now show that $\tilde{Q}$ is an intersection support for $(G,\mathcal{H,K})$. Let $K\in\mathcal{K}$ be arbitrary. We claim that $\tilde{Q}[\mathcal{H}_{K}]$ induces a connected subgraph of $\tilde{Q}$. Since $\tilde{Q}^{\prime}$ is an intersection support for $(G,\mathcal{H}^{*},\mathcal{K})$, the induced subgraph $\tilde{Q}^{\prime}[\mathcal{H}^{*}_{K}]$ of $\tilde{Q}^{\prime}$ is connected. By definition, $\mathcal{H}^{*}\subseteq\mathcal{H}$ consists of precisely the maximal elements of $(\mathcal{H},\preceq_{N})$. Therefore, for any $H\in\mathcal{H}_{K}\setminus\mathcal{H}^{*}_{K}$, $\exists H^{\prime}\in\mathcal{H}^{*}$ such that $H\subseteq H^{\prime}$ and $\{H,H^{\prime}\}$ is an edge in $\tilde{Q}$. But $H^{\prime}$ intersects $K$ since $H$ interacts $K$. It follows that $H^{\prime}\in\mathcal{H}^{*}_{K}$ and hence, the induced subgraph $\tilde{Q}[\mathcal{H}_{K}]$ of $\tilde{Q}$ is connected. This completes the proof.

### v3

@@ Abstract
All our results are algorithmic and yield polynomial time algorithms (when the treewidth is bounded). The existence and construction of sparse supports is a crucial step in the design and analysis of PTAS for several packing and covering problems.
@@ 1 Introduction
A hypergraph is defined by a set $X$ of vertices, and a collection $\mathcal{H}$ of subsets of $X$ called hyperedges. A support for $(X,\mathcal{H})$ is a graph $Q=(X,F)$ s.t. $Q[H]$ is connected for each $H\in\mathcal{H}$, where $Q[H]$ is the subgraph of $Q$ induced by the vertices in $H$. A complete graph on $X$ is clearly a support, but the existence of a support is non-trivial if we restrict the support to be a sparse graph. For example, Johnson and Pollack [Johnson1987Hypergraph] showed that it is NP-complete to decide if a hypergraph admits a planar support. Buchin et al. [buchin2011planar] showed that it is NP-hard to decide if s hypergraph admits a support that is k-outerplanar for any $k\geq 2$.
Voloshina and Feinberg introduced planar supports as a notion of planarity of hypergraphs [voloshina1984planarity]. Since, then the existence and construction of planar supports has found applications in several domains such as visualization [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored] and network design [anceaume2006semantic, baldoni2007tera, baldoni2007efficient, chand2005semantic, hosoda2012approximability, korach2003clustering, onus2011minimum]. The existence of sparse supports and similar questions have also found applicability for geometric problems [Mustafa17, BasuRoy2018, Cohen-AddadM15, krohn2014guarding, mustafa2010improved, RR18].
In this paper, we consider hypergraphs defined on a host graph. That is, we are given as input a graph $G=(V,E)$ along with a collection $\mathcal{H}$ of subgraphs of $G$, and our goal is to construct a support for a hypergraph defined on $G$. We focus on restrictions on $G$ and the subgraphs $\mathcal{H}$ that guarantee the existence of a sparse support.
Our motivation comes from the fact that the existence of a support implies a PTAS for a large class of packing and covering problems via a now well-established technique. See for example, [DBLP:conf/walcom/AschnerKMY13, ChanH12, mustafa2010improved, RR18, raman2025supportsgraphsboundedgenus]. We now describe our general framework before giving some potential applications.
Let $G=(V,E)$ be a given graph and let $\mathbf{b}(V)\subseteq V$ be a set of terminals. Let $\mathbf{r}(V)$ denote the vertices in $V\setminus\mathbf{b}(V)$. The set $\mathbf{r}(V)$ constitute the non-terminals. A family $\mathcal{H}$ of subgraphs of $G$ defines a hypergraph $(\mathbf{b}(V),\mathcal{H})$, whose vertex set is $\mathbf{b}(V)$ and each $H\in\mathcal{H}$ defines a hyperedge consisting of the vertices in $\mathbf{b}(V)\cap V(H)$. We call this hypergraph the primal hypergraph. Similarly, we define the dual hypergraph whose elements are the subgraphs in $\mathcal{H}$, and each $v\in\mathbf{b}(V)$ defines a hyperedge consisting of all $H\in\mathcal{H}$ containing $v$. We also consider a generalization of the primal and dual hypergraphs namely, the intersection hypergraphs. Let $\mathcal{H}$ and $\mathcal{K}$ be two families of subgraphs of $G$. The intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ is the hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}}),\mbox{ where }\mathcal{H}_{K}=\{H\in\mathcal{H}:V(H)\cap V(K)\neq\emptyset\}$
@@
A support for a primal (resp. dual/intersection) hypergraph is called a primal (resp. dual/intersection) support. Formally, a primal support is a graph $Q$ on $\mathbf{b}(V)$ such that for each $H\in\mathcal{H}$, the vertices $V(H)\cap\mathbf{b}(V)$ induce a connected subgraph of $Q$. A dual support is a graph $Q^{*}$ on $\mathcal{H}$ such that for each $v\in\mathbf{b}(V)$, the elements in $\{H\in\mathcal{H}:v\in H\}$ induce a connected subgraph of $Q^{*}$. Finally, an intersection support is a graph $\tilde{Q}$ on $\mathcal{H}$ such that for each $K\in\mathcal{K}$, the elements in $\mathcal{H}_{K}$ induce a connected subgraph of $\tilde{Q}$. In the dual setting, note that we can assume without loss of generality that $\mathbf{b}(V)=V$.
See Figure 2 in [raman2025supportsgraphsboundedgenus] for examples of primal, dual and intersection hypergraphs arising from a given host graph, and their corresponding primal, dual, and intersection supports.
We design algorithms to construct primal, dual and intersection supports for graphs of bounded treewidth. We obtain similar results for a special case of graphs of treewidth 2, namely, the outerplanar graphs.
@@
{problem}
{problem}
@@
As we show below, for the problem defined above, the answer is negative even if $\mathcal{G}$ is the family of trees. However, if we restrict the family of subgraphs to be non-piercing, then the answer to the question above is affirmative. A family $\mathcal{H}$ of connected subgraphs of a graph $G$ is said to be non-piercing, if for each $H,H^{\prime}\in\mathcal{H}$, both $V(H)\setminus V(H^{\prime})$ and $V(H^{\prime})\setminus V(H)$ induce connected subgraphs of $G$.
The motivation to consider hypergraphs defined by non-piercing subgraphs of a graph comes from geometry. Raman and Ray [RR18] showed that the intersection hypergraph of non-piercing regions [^2] A family $\mathcal{R}$ of connected regions in $\mathbb{R}^{2}$ is non-piercing if for each $R,R^{\prime}\in\mathcal{R}$, both $R\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are connected regions. They generalize disks, pseudodisks, and halfspaces in $\mathbb{R}^{2}$ in the plane admits a planar support. It was shown in [raman2025supportsgraphsboundedgenus] that non-piercing regions in $\mathbb{R}^{2}$ correspond to non-piercing subgraphs of a planar graph. The work was extended to surfaces of higher genus where the subgraphs satisfy a condition of being cross-free. The authors showed that the intersection hypergraph defined by cross-free subgraphs of a bounded genus graph admits a support of bounded genus. It was observed that non-piercing is an insufficient condition to obtain a dual support of bounded genus. In this paper, we show that the non-piercing condition is sufficient to obtain an intersection support of bounded treewidth. In order to construct an intersection support, we require the construction of primal and dual supports.
@@
Since the graph classes we consider here are simple, one may wonder if the non-piercing condition is necessary to obtain sparse supports. The following examples show that this is indeed the case. For the primal hypergraph, consider a star $K_{1,n}$ with leaves $v_{1},\ldots,v_{n}$, and $v_{0}$, the central vertex. We color the leaves $\mathbf{b}$ and the central vertex $\mathbf{r}$. The subgraphs in $\mathcal{H}$ consist of all pairs of leaves along with the central vertex $v_{0}$. Thus, =$\mathcal{H}=\{H_{ij}:H_{ij}=\{v_{i},v_{0},v_{j}\}\}$. It is easy to see that the primal support in this case is $K_{n}$, the complete graph on $n$ vertices. For the dual hypergraph, consider a star $K_{1,\binom{n}{2}}$ with central vertex $v_{0}$. Each leaf $v_{j}$ is labeled by a unique pair $\{x_{j},y_{j}\}$ of $\{1,\ldots,n\}$. There are $n$ subgraphs $H_{1},H_{2},\ldots,H_{n}$, where $H_{i}=\{v_{0}\}\cup\{v_{j}:i\in\{x_{j},y_{j}\}\}$. It is easy to check that the dual support in this case is also $K_{n}$, as there are exactly two subgraphs containing each leaf.
Below, we give a few potential applications where our techniques can apply. Since our focus is on the existence and construction of supports, we do not elaborate more on the applications as the algorithmic results follow via standard techniques once there is a support graph.
Consider a graph $G$. Note that a set of cliques in $G$ is a natural non-piercing family of subgraphs of $G$. Given two sets of cliques of $G$, a set of red cliques and a set of blue cliques, consider the problem of choosing the smallest number of red cliques s.t. each blue clique shares a vertex with at least one of the chosen red cliques. If $G$ has bounded treewidth, since we show that there exists an intersection support of bounded treewidth, it implies that this problem admits a PTAS via a standard local-search framework described in [DBLP:conf/walcom/AschnerKMY13, ChanH12, mustafa2010improved]. Note that in this problem, the set of red/blue cliques can be specified arbitrarily.
As another potential example, let $\mathcal{C}$ be a set of cliques of a graph $G$. Then, the problem of choosing the largest subset $S$ of $\mathbf{b}(V)$ so that each clique in $\mathcal{C}$ contains at most one vertex of $S$ admits a sub-exponential time algorithm for graphs with treewidth at most $c\log n$, for some sufficiently small $c<1$. The sub-exponential time algorithm is as follows: we show that a support graph $Q$ can be computed in polynomial time, and $Q$ has treewidth at most $2^{O(\textsc{tw}(G))}=n^{\delta}$ for some $\delta<1$. We enumerate over the subsets of vertices in a balanced separator of $Q$ and then recursively find such subgraphs in the two sub-problems obtained on removing the separator while ensuring that the choices made in the sub-problems are compatible with the choices made further up the recursion tree. Since $Q$ is a support, each clique that contains vertices on either side of the separator must necessarily contain vertices in the separator. Since the separator is balanced, the recursion tree has depth $O(\log n)$ and since the number of vertices of $Q$ is at most $n$, and the separator has sublinear size, we obtain a sub-exponential time algorithm.
@@ 2 Related Work
The notion of planarity for hypergraphs, unlike that for graphs is not uniquely defined. This notion was first studied by Zykov [zykov], but that was very restrictive. Voloshina and Feinberg [voloshina1984planarity] defined a notion of planarity of hypergraphs that is equivalent to the existence of a support graph that is planar, i.e., a planar support. Johnson and Pollak [Johnson1987Hypergraph] showed that deciding if a hypergraph admits a planar support is NP-hard. Buchin et al. [buchin2011planar] strengthened their result showing that it is NP-hard even to decide if a hypergraph admits a support that is $k$-outerplanar, for $k\geq 2$. However, one can decide in polynomial time if a hypergraph has a support that is a path, a cycle, a cactus [^3] A connected graph is called cactus if each of its edges participates in at most one cycle., or a tree with bounded maximum degree [beeri, BOOTH1976335, brandes2010blocks, buchin2011planar, tarjan1984simple]. Recently, Chudnovsky et al. [chudnovsky2025vertex] gave characterizations of hypergraphs having a tree support.
@@ 3 Preliminaries
Let $\mathbf{b}(V)$ and $\mathbf{r}(V)$ be a partition of $V(G)$ into a set of terminals and a set of non-terminals, respectively. Formally, let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be a 2-coloring of $V(G)$ (not necessarily proper). We call $\mathbf{r}(V)=c^{-1}(\mathbf{r})$ the set of red vertices and $\mathbf{b}(V)=c^{-1}(\mathbf{b})$ the set of blue vertices. We use $\mathbf{b}(H)$ to denote $V(H)\cap\mathbf{b}(V)$, and $\mathbf{r}(H)$ to denote $V(H)\cap\mathbf{r}(V)$.
Definition 1 (Non-piercing subgraphs).
@@
For a graph $G$ and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs of $G$, we use the pair $(G,\mathcal{H})$ to refer to the primal or dual hypergraph, and triple $(G,\mathcal{H},\mathcal{K})$ for the intersection hypergraph. We call the tuple $(G,\mathcal{H})$ a graph system and the tuple $(G,\mathcal{H},\mathcal{K})$ an intersection system. If $G\in\mathcal{G}$, and $\mathcal{H,K}$ satisfy a property $\Pi$, then we say that $(G,\mathcal{H})$ is a $\mathcal{G}$ $\Pi$ graph system, or that $(G,\mathcal{H},\mathcal{K})$ is a $\mathcal{G}$ $\Pi$ intersection system. For example, if $G$ has bounded treewidth, and $\mathcal{H}$ and $\mathcal{K}$ are connected subgraphs of $G$, then $(G,\mathcal{H},\mathcal{K})$ is a bounded treewidth connected intersection system. Similarly, if $G$ is outerplanar and $\mathcal{H}$ is non-piercing, then $(G,\mathcal{H})$ is an outerplanar non-piercing system.
Note 5.
Since two subgraphs are said to intersect if they share a vertex, a hypergraph defined by subgraphs of a graph $G$ remains the same if the subgraphs are replaced by the corresponding induced subgraphs. Moreover, if $\mathcal{H}$ is a collection of non-piercing subgraphs, then it remains non-piercing if the subgraphs are replaced by the corresponding induced subgraphs. Therefore, we assume throughout that the term subgraph refers to an induced subgraph of $G$.
Also note that adding edges to $G$ leaves the subgraphs in $\mathcal{H}$ or $\mathcal{K}$ non-piercing. Thus, we can assume in the outerplanar setting that $G$ is a maximal outerplanar graph, and in the setting where $G$ has bounded treewidth, we can assume that each bag in a tree decomposition of the graph induces a complete graph, as this does not increase the treewidth (see Section 5 for more details).
@@
We now show that $\tilde{Q}$ is an intersection support for $(G,\mathcal{H,K})$. Let $K\in\mathcal{K}$ be arbitrary. We claim that $\tilde{Q}[\mathcal{H}_{K}]$ induces a connected subgraph of $\tilde{Q}$. Since $\tilde{Q}^{\prime}$ is an intersection support for $(G,\mathcal{H}^{*},\mathcal{K})$, the induced subgraph $\tilde{Q}^{\prime}[\mathcal{H}^{*}_{K}]$ of $\tilde{Q}^{\prime}$ is connected. By definition, $\mathcal{H}^{*}\subseteq\mathcal{H}$ consists of precisely the maximal elements of $(\mathcal{H},\preceq_{N})$. Therefore, for any $H\in\mathcal{H}_{K}\setminus\mathcal{H}^{*}_{K}$, $\exists H^{\prime}\in\mathcal{H}^{*}$ such that $H\subseteq H^{\prime}$ and $\{H,H^{\prime}\}$ is an edge in $\tilde{Q}$. But $H^{\prime}$ intersects $K$ since $H$ interacts $K$. It follows that $H^{\prime}\in\mathcal{H}^{*}_{K}$ and hence, the induced subgraph $\tilde{Q}[\mathcal{H}_{K}]$ of $\tilde{Q}$ is connected. This completes the proof. ∎